OpenStudy (anonymous):

Which decimal is equivelent to 8 9/12

OpenStudy (johnweldon1993):

Well first we should turn this mixed number into an improper fraction....you do this by multiplying the denominator (12) by the whole number (8) and then adding the numerator... so 12 * 8 = 96... 96 + 9 = 105 And you put this number over the original denominator (12) so \[\large \frac{105}{12}\] Now just plug this into a calculator...

OpenStudy (anonymous):

Ok haha that is confusing:/

OpenStudy (johnweldon1993):

Hmm...well this particular one could be a little easier...but I showed you that way because it wont always be the case... \[\large 8\frac{9}{12}\] We can simplify 9/12 a bit right? this is the same as 3/4 because both are divisable by 3...so now we have \[\large 8\frac{3}{4}\] We know that 3/4 = .75 right? well then we have 8.75

OpenStudy (johnweldon1993):

but like I said...this won't always be the case...what is this was \[\large 8\frac{11}{12}\] Not so much fun...but regardless You can always evaluate the smaller fraction first....so 11/12 = .916666 and then this would just equal 8.916666 etc... a little better?